OpenStudy (1645323):

The profit for a certain company can be modeled by P(t) = t3 + 5t2 - 4t - 20, where t is the time in days. Use factoring to determine the values of t for which P(t) = 0.

OpenStudy (cwrw238):

as the last term is 20 the roots of the equation will be factors of 20 us the factor theorem to find one root P(2) = 8 +20 - 8 - 20 = 0 so t - 2 is one factor now you can use long division or synthetic division to find the other factors

OpenStudy (1645323):

the other one is t-5?

OpenStudy (1645323):

i mean 4

OpenStudy (cwrw238):

sorry i was wrong to say all the roots would be factors of 20 2 certainly is but the otrhers may not be whole numbers

OpenStudy (cwrw238):

its best to do the division at this stage

OpenStudy (cwrw238):

the quotient is t^2 + 7t +10 can you factor this?

OpenStudy (cwrw238):

do you agree with that result ?

OpenStudy (1645323):

yes (t+5)(t+2)

OpenStudy (cwrw238):

ok so now you have your values of t from P(t) = (t -2)(t + 2)(t + 5) = 0

OpenStudy (1645323):

and those are all the roots?

OpenStudy (cwrw238):

roots are 2, -5 and -2

OpenStudy (1645323):

ok thansk

OpenStudy (cwrw238):

yw